The length of life of an instrument produced

by a machine has a normal distribution with a

mean of 12 months and standard deviation of

2 months. Find the probability that an

instrument produced by this machine will last

a) less than 7 months.

b) between 7 and 12 months.


1
Expert's answer
2021-10-18T13:38:00-0400

a)

"z=\frac{x-\mu}{\sigma}=\frac{7-12}{2}=-2.5"


"P(x<7)=(z<-2.5)=0.0062"


b)

"z_1=\frac{x_1-\mu}{\sigma}=\frac{7-12}{2}=-2.5"


"z_2=\frac{x_2-\mu}{\sigma}=\frac{12-12}{2}=0"


"P(7<x<12)=P(z_1<z<z_2)="

"=P(x<12)-P(x<7)=0.5-0.0062=0.4938"